# Spectrum of Schrödinger operators on subcovering graphs

### v1

@@ 4.2. Subcoverings of the $d$-dimensional lattice
i) Let $\varrho_{s}:=t_{s1}+\ldots+t_{sd}$ be even for all $s\in{\mathbb{N}}_{d_{o}}$. Since $\pi 1\!\!1_{d}\in K_{1}^{+}$, due to Proposition 2.4.ii, we have $\lambda_{1}^{+}(\mathfrak{t})=\lambda_{1}^{+}=4d$.
@@ 4.3. Subcoverings of the diamond lattice
between the vertex $v_{1}+\mathfrak{a}$ in a lower copy and the vertex $v_{2}+\mathfrak{a}_{3}+\mathfrak{a}$ in the next copy, $\mathfrak{a}\in\Gamma$, where $v_{1},v_{2}$ are the vertices of ${\bf G}$ from the fundamental cell $\Omega$, and $\Gamma$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$. The diamond lattice ${\mathcal{G}}$ is a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}$ consisting of two vertices $v_{1},v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d.
First we describe the spectrum of the Schrödinger operator $H=\Delta+Q$ on the diamond lattice ${\mathcal{G}}$ with a $\Gamma$-periodic potential $Q$ defined by (2.13).
Lemma 4.4.
The spectrum of the Schrödinger operator $H=\Delta+Q$ on the diamond lattice ${\mathcal{G}}$ with the potential $Q$ defined by (2.13) is given by
@@
The set $K_{1}^{+}=K_{2}^{-}$ defined by (4.9) is an algebraic curve of co-dimension 2 shown in red line in Fig. 6. The band edges $\lambda_{1}^{+}$ and $\lambda_{2}^{-}$ are degenerate. This example of periodic graphs where the non-degeneracy assumption fails was constructed in [BK20].
Proof. For the diamond lattice ${\mathcal{G}}$, the matrix of the Floquet operator $H(k)$ defined by (1.4) in the standard basis of $\ell^{2}(\{v_{1},v_{2}\})\cong{\mathbb{C}}^{2}$ has the form
$H(k)=\left(\begin{array}[]{cc}4+q&-1-e^{ik_{1}}-e^{ik_{2}}-e^{ik_{3}}\\ -1-e^{-ik_{1}}-e^{-ik_{2}}-e^{-ik_{3}}&4-q\end{array}\right),$
$k=(k_{1},k_{2},k_{3})\in{\mathcal{B}}=(-\pi,\pi]^{3}$. The eigenvalues of $H(k)$ are given by
$\lambda_{j}(k)=4+(-1)^{j}\sqrt{f(k)+q^{2}},\qquad j=1,2,\quad$
$\textrm{where}\qquad f(k)=|1+e^{ik_{1}}+e^{ik_{2}}+e^{ik_{3}}|^{2}.$
We have
$\min\limits_{k\in{\mathcal{B}}}f(k)=f\big{(}K_{1}^{+}\big{)}=0,\qquad\max\limits_{k\in{\mathcal{B}}}f(k)=f(0,0,0)=16,$
where $K_{1}^{+}$ is defined by (4.9). Then, using the definition of the band edges $\lambda_{j}^{\pm}$ in (1.6), we obtain that the spectrum of $H$ has the form (4.8), (4.9).  $\Box$
@@
Then $k_{o}\in K_{1}^{+}=K_{2}^{-}$, where $K_{1}^{+}=K_{2}^{-}$ is defined by (4.9) and for $k_{o}$ and ${\mathcal{T}}=(t_{1},t_{2},t_{3})$ the condition (2.5) is fulfilled. For the point $0\in K_{1}^{-}=K_{2}^{+}$ the condition (2.5) is also fulfilled. Thus, due to Corollary 2.5, the diamond lattice ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, i.e., $\sigma(H)=\sigma(H_{\mathfrak{t}})$.
Now let $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ be a primitive set of the lattice $\Gamma$, where $\mathfrak{t}_{s}=t_{s1}\mathfrak{a}_{1}+t_{s2}\mathfrak{a}_{2}+t_{s3}\mathfrak{a}_{3}$, $s=1,2$. If for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$, then using the same arguments as above we obtain that ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral. Conversely, if ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral, then, due to Corollary 2.5,

### v2

@@ 4.2. Subcoverings of the $d$-dimensional lattice
i) Let $\varrho_{s}:=t_{s1}+\ldots+t_{sd}$ be even for all $s\in{\mathbb{N}}_{d_{o}}$. Since $\pi 1\!\!1_{d}\in K_{1}^{+}$, due to Proposition 2.2.i, we have $\lambda_{1}^{+}(\mathfrak{t})=\lambda_{1}^{+}=4d$.
@@ 4.3. Subcoverings of the diamond lattice
between the vertex $v_{1}+\mathfrak{a}$ in a lower copy and the vertex $v_{2}+\mathfrak{a}_{3}+\mathfrak{a}$ in the next copy of ${\bf G}$, $\mathfrak{a}\in\Gamma$, where $v_{1},v_{2}$ are the vertices of ${\bf G}$ from the fundamental cell $\Omega$, and $\Gamma$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$. The diamond lattice ${\mathcal{G}}$ is a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}$ consisting of two vertices $v_{1},v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d.
The spectrum of the Schrödinger operator $H=\Delta+Q$ on the diamond lattice ${\mathcal{G}}$ with a $\Gamma$-periodic potential $Q$ defined by (2.12) is given by
$\sigma(H)=[\lambda_{1}^{-},\lambda_{1}^{+}]\cup[\lambda_{2}^{-},\lambda_{2}^{+}],$
@@
The set $K_{1}^{+}=K_{2}^{-}$ defined by (4.7) is shown in red line in Fig. 6. The band edges $\lambda_{1}^{+}$ and $\lambda_{2}^{-}$ are degenerate. This example of periodic graphs where the non-degeneracy assumption fails was constructed in [BK20].
@@
Then $k_{o}\in K_{1}^{+}=K_{2}^{-}$, where $K_{1}^{+}=K_{2}^{-}$ is defined by (4.7) and for $k_{o}$ and ${\mathcal{T}}=(t_{1},t_{2},t_{3})$ the condition (2.3) is fulfilled. For the point $0\in K_{1}^{-}=K_{2}^{+}$ the condition (2.3) is also fulfilled. Thus, due to Proposition 2.2.i, the diamond lattice ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, i.e., $\sigma(H)=\sigma(H_{\mathfrak{t}})$.
Now let $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ be a primitive set of the lattice $\Gamma$, where $\mathfrak{t}_{s}=t_{s1}\mathfrak{a}_{1}+t_{s2}\mathfrak{a}_{2}+t_{s3}\mathfrak{a}_{3}$, $s=1,2$. If for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$, then using the same arguments as above we obtain that ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral. Conversely, if ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral, then, due to Proposition 2.2.i,
@@
Acknowledgments. The work of the author was supported by the RSF grant No. 25-21-00157.
Data availability No datasets were generated or analysed during the current study.
Conflict of interest None declared.